# Li He1, Pei-Yang Lin1, Peng-Hu Wei5, Guo-Guang Zhao5, Yun-Zhe Liu†,1,4 1

### v1

@@ C.1 Du-IN VQ-VAE
The architecture of the Du-IN VQ-VAE model contains three parts: (1) Neural Tokenizer, (2) Vector Quantizer, and (3) Neural Regressor. The overall architecture of "Neural Tokenizer" is shown in Figure 2. The "Vector Quantizer" is implemented similarly in LaBraM[19]. The "Neural Regressor" contains:
@@
Neural Tokenizer (Neural Transformer) Spatial Encoder Linear Projection $10\rightarrow 16$
@@
Embedding-to-Codex Projection $160\rightarrow 160(Tanh)\rightarrow 64$
@@
Neural Regressor Transformer Decoder # of Transformer Layers 4
@@ C.2 Du-IN MAE
The architecture of the Du-IN MAE model contains two parts: (1) Neural Encoder, and (2) Token Prediction Head. The overall architecture of the "Neural Encoder" is shown in Figure 2. The hyperparameters of "Neural Encoder" are the same as those of "Neural Tokenizer" in Du-IN VQ-VAE. It’s worth noting that when training Du-IN MAE, the weights of the "Neural Encoder" are randomly initialized, instead of loaded from the pre-trained Du-IN VQ-VAE model. The hyperparameters for Du-IN MAE training are shown in Table 6.
@@ C.3 Du-IN CLS
The architecture of the Du-IN CLS model contains two parts: (1) Neural Encoder, and (2) Label Prediction Head. The overall architecture of the "Neural Encoder" is shown in Figure 2. The hyperparameters of "Neural Encoder" are the same as those of "Neural Tokenizer" in Du-IN VQ-VAE. It’s worth noting that the "Neural Encoder" weights in Du-IN CLS can be loaded from either the pre-trained Du-IN MAE or the pre-trained Du-IN VQ-VAE. In the ablation experiments shown in Table 2, our models have different suffixes:
@@
  • Du-IN (vqvae+vq): The weights of the "Neural Encoder" in the Du-IN CLS model are loaded from the pre-trained Du-IN VQ-VAE. When fine-tuning it on the downstream task, the "Vector Quantizer" in the pre-trained Du-IN VQ-VAE is inserted between "Neural Encoder" and "Label Prediction Head". This is the same operation in DeWave[15].
  • Du-IN (vqvae): The weights of the "Neural Encoder" in the Du-IN CLS model are loaded from the pre-trained Du-IN VQ-VAE. This is the same operation in EEGFormer [9].
  • Du-IN (mae): The weights of the "Neural Encoder" in the Du-IN CLS model are loaded from the pre-trained Du-IN MAE. This is the same operation in LaBraM [19].
@@
Linear Projection $30\times 160\rightarrow 128(ReLU)\rightarrow 61$
@@ Appendix E Du-IN Pre-training Analysis
where (1) $q_{\phi}(z|x)$ denotes the neural tokenizer that obtains neural tokens; (2) $p_{\psi}(x|z)$ decodes the original sEEG signal given input neural tokens; (3) $p_{\theta}(z|\tilde{x})$ recovers the neural tokens based on the masked sEEG signal, which is our Du-IN pre-training task.
The whole framework is optimized through a two-stage procedure as [36, 30]. For the first stage, we train the neural tokenizer as a discrete variational autoencoder by minimizing the reconstruction loss $-\mathbb{E}_{z_{i}\sim q_{\phi}(\bm{\mathrm{z}}|x_{i})}\mathrm{log}\ p_{\psi}(\tilde{x}_{i}|z_{i})$ with a uniform prior. For the second stage, we set $q_{\phi}$ as well as $p_{\psi}$ fixed and learn the prior $p_{\theta}$ by minimizing the loss $D_{\mathrm{KL}}$. For simplicity, $q_{\phi}(\bm{\mathrm{z}}|x_{i})$ is defined as a one-point distribution with the most likely neural tokens $\hat{z}_{i}=\mathop{\arg\max}\limits_{z}q_{\phi}(\bm{\mathrm{z}}|x_{i})$. Consequently, we can rewrite Equation 10 as
@@ Appendix F Visualization of Vector-Quantized sEEG Regression
We further visualize how the sEEG signals are reconstructed. As depicted in Figure 7, although some details are missing, the overall trend of the signals is reconstructed well. Meanwhile, there is a stable decrease in the reconstruction loss during training, which indicates the discrete codebook does learn high-level information from sEEG signals.
@@ Appendix H Channel Contribution Analysis
Appendix I Subject-Wise Evaluation
@@
Methods Pre-trained Accuracy (%) $\pm$ Std (%)
subj-01 subj-02 subj-03 subj-04
EEG-Conformer[34] 58.41$\pm$1.03 69.82$\pm$1.22 19.50$\pm$1.71 49.65$\pm$2.38
CNN-BiGRU[27] 46.46$\pm$4.03 68.06$\pm$1.56 4.35$\pm$0.44 17.68$\pm$3.88
@@
Methods Pre-trained Accuracy (%) $\pm$ Std (%)
subj-05 subj-06 subj-07 subj-08
EEG-Conformer[34] 65.44$\pm$1.31 31.06$\pm$2.58 47.89$\pm$1.86 42.12$\pm$2.08
CNN-BiGRU[27] 51.26$\pm$4.93 31.52$\pm$1.48 47.75$\pm$1.12 24.64$\pm$4.44
@@
Methods Pre-trained Accuracy (%) $\pm$ Std (%)
subj-09 subj-10 subj-11 subj-12
EEG-Conformer[34] 56.51$\pm$1.98 22.22$\pm$1.07 57.10$\pm$2.03 29.87$\pm$1.44
CNN-BiGRU[27] 44.03$\pm$5.88 7.11$\pm$0.71 28.44$\pm$3.42 13.17$\pm$3.41
@@
Appendix J Subject-Wise Electrode Locations

### v3

@@ C.1 Du-IN VQ-VAE
The architecture of the Du-IN VQ-VAE model contains three parts: (1) Du-IN Encoder, (2) Vector Quantizer, and (3) Du-IN Regressor. The overall architecture of "Du-IN Encoder" is shown in Figure 2. The "Vector Quantizer" is implemented similarly in LaBraM[28]. The "Du-IN Regressor" contains:
@@
Du-IN Encoder Spatial Encoder Linear Projection $10\rightarrow 16$
@@
Embedding-to-Codex Projection $160\rightarrow 160(\mathrm{Tanh})\rightarrow 64$
@@
Du-IN Regressor Transformer Decoder # of Transformer Layers 4
@@ C.2 Du-IN MAE
The architecture of the Du-IN MAE model contains two parts: (1) Du-IN Encoder, and (2) Token Prediction Head. The overall architecture of the "Du-IN Encoder" is shown in Figure 2. The hyperparameters of "Du-IN Encoder" are the same as those in Du-IN VQ-VAE. It’s worth noting that when training Du-IN MAE, the weights of the "Du-IN Encoder" are randomly initialized, instead of loaded from the pre-trained Du-IN VQ-VAE model. The hyperparameters for Du-IN MAE training are shown in Table 6.
@@ C.3 Du-IN CLS
The architecture of the Du-IN CLS model contains two parts: (1) Du-IN Encoder, and (2) Label Prediction Head. The overall architecture of the "Du-IN Encoder" is shown in Figure 2. The hyperparameters of "Du-IN Encoder" are the same as those in Du-IN VQ-VAE. It’s worth noting that the "Du-IN Encoder" weights in Du-IN CLS can be loaded from either the pre-trained Du-IN MAE or the pre-trained Du-IN VQ-VAE. In the ablation experiments shown in Table 2, our models have different suffixes:
@@
  • Du-IN (vqvae+vq): The weights of the "Du-IN Encoder" in the Du-IN CLS model are loaded from the pre-trained Du-IN VQ-VAE. When fine-tuning it on the downstream task, the "Vector Quantizer" in the pre-trained Du-IN VQ-VAE is inserted between "Du-IN Encoder" and "Label Prediction Head". This is the same operation in DeWave[19].
  • Du-IN (vqvae): The weights of the "Du-IN Encoder" in the Du-IN CLS model are loaded from the pre-trained Du-IN VQ-VAE. This is the same operation in EEGFormer [12].
  • Du-IN (mae): The weights of the "Du-IN Encoder" in the Du-IN CLS model are loaded from the pre-trained Du-IN MAE. This is the same operation in LaBraM [28].
  • Du-IN (poms): The weights of the "Du-IN Encoder" in the Du-IN CLS model are loaded from the Du-IN MAE, which is pre-trained on multiple subjects. The modification of the Du-IN VQ-VAE and the Du-IN MAE to support multi-subject pre-training includes (1) initializing different spatial encoders for different subjects and (2) sharing the same transformer encoder and neural codex.
@@
Linear Projection $30\times 160\rightarrow 128(\mathrm{ReLU})\rightarrow 61$
@@ Appendix E Du-IN Pre-training Analysis
where (1) $q_{\phi}(z|x)$ denotes the Du-IN Encoder in the Du-IN VQ-VAE that obtains neural tokens; (2) $p_{\psi}(x|z)$ decodes the original sEEG signal given input neural tokens; (3) $p_{\theta}(z|\tilde{x})$ recovers the neural tokens based on the masked sEEG signal, which is our Du-IN pre-training task.
The whole framework is optimized through a two-stage procedure as [47, 40]. For the first stage, we train the Du-IN Encoder in the Du-IN VQ-VAE as a discrete variational autoencoder by minimizing the reconstruction loss $-\mathbb{E}_{z_{i}\sim q_{\phi}(\bm{\mathrm{z}}|x_{i})}\mathrm{log}\ p_{\psi}(\tilde{x}_{i}|z_{i})$ with a uniform prior. For the second stage, we set $q_{\phi}$ as well as $p_{\psi}$ fixed and learn the prior $p_{\theta}$ by minimizing the loss $D_{\mathrm{KL}}$. For simplicity, $q_{\phi}(\bm{\mathrm{z}}|x_{i})$ is defined as a one-point distribution with the most likely neural tokens $\hat{z}_{i}=\mathop{\arg\max}\limits_{z}q_{\phi}(\bm{\mathrm{z}}|x_{i})$. Consequently, we can rewrite Equation 10 as
@@ Appendix F Visualization of Vector-Quantized sEEG Regression
We further visualize how the sEEG signals are reconstructed. As depicted in Figure 7, although some details are missing, the overall trend of the signals is reconstructed well. Meanwhile, there is a stable decrease in the reconstruction loss during training, which indicates the discrete codex does learn high-level information from sEEG signals.
@@ Appendix H Channel Contribution Analysis
Appendix I Effectiveness of Region-Specific Channel Selection
DeWave [19] successfully reconstructs 128-channel EEG signals with the same setting of vector-quantizer. However, this is not the case under the sEEG setting, which is shown in Table 8. It’s worth noting that sEEG signals are fundamentally different from EEG signals due to (1) the high information density and (2) the high specificity of different regions. Due to the desynchronization nature [7] of the brain during awake tasks, only specific brain regions are related to tasks. Therefore, only after region-specific channel selection, the Du-IN VQ-VAE model can successfully reconstruct the original signals, thus identifying the fine-grained state of brain regions.
Table 8: Ablations to validate the effectiveness of region-specific channel selection.
       Settings        Setting 1        Setting 2        Setting 3
       MSE        0.2969$\pm$0.0376        0.5211$\pm$0.0492        0.9673$\pm$0.0148
  • 1
    Setting 1: Select top-10 channels relevant to speech decoding for neural signal reconstruction.
  • 2
    Setting 2: Randomly select 10 channels for neural signal reconstruction.
  • 3
    Setting 3: Use all channels (109.75 channels on average) for neural signal reconstruction.
Appendix J Additional Group-Wise Evaluation
The cross-entropy loss of different methods from each subject is provided in Table 9, with the best in bold and the second underlined. For model comparison, we report the average and standard deviation values (within each subject) on six different random seeds to obtain comparable results. "Std" means standard deviation.
Table 9: The cross-entropy loss of different methods (with the best in bold and the second underlined).
Methods Token Level Config Model Size Cross-Entropy $\pm$ Ste
PT11footnotemark: 1 MS22footnotemark: 2
TS-TCC[20] Region 0.32M 3.8871$\pm$0.3072
CNN-BiGRU[37] Region - 0.54M 4.0294$\pm$0.7621
EEG-Conformer[44] Region - 2.34M 3.8165$\pm$0.3456
Neuro-BERT[50] Region 2.14M 3.6416$\pm$0.4360
DeWave[19] Region - 5.70M 4.1891$\pm$0.5722
BrainBERT[49] Channel 43.58M 4.6254$\pm$0.1984
BrainBERT[49] Channel 43.58M 4.6190$\pm$0.2132
Brant[54] Channel 69.35M 4.7962$\pm$0.7082
Brant[54] Channel 69.35M 5.0294$\pm$1.0621
LaBraM[28] Channel 6.85M 4.8591$\pm$0.2723
LaBraM-PopT[28, 10] Channel 6.85M 4.6564$\pm$0.1893
Du-IN Region - 4.38M 3.5083$\pm$0.3003
Du-IN (vqvae+vq) Region 4.38M 3.7244$\pm$0.3104
Du-IN (vqvae) Region 4.38M 3.4309$\pm$0.2781
Du-IN (mae) Region 4.38M 3.3707$\pm$0.2882
Du-IN (poms) Region 5.18M 3.4429$\pm$0.2754
  • 1
    PT: Whether the model is pre-trained before evaluation.
  • 2
    MS: Whether the model is pre-trained across multiple subjects.
Appendix K Subject-Wise Evaluation
@@ Appendix I Effectiveness of Region-Specific Channel Selection
Methods Config Accuracy (%) $\pm$ Std (%)
PT11footnotemark: 1 MS22footnotemark: 2 subj-01 subj-02 subj-03 subj-04
TS-TCC[20] 26.90$\pm$1.72 61.57$\pm$1.21 6.65$\pm$1.09 20.66$\pm$0.79
CNN-BiGRU[37] - 46.46$\pm$4.03 68.06$\pm$1.56 4.35$\pm$0.44 17.68$\pm$3.88
EEG-Conformer[44] - 58.41$\pm$1.03 69.82$\pm$1.22 19.50$\pm$1.71 49.65$\pm$2.38
Neuro-BERT[50] 60.44$\pm$2.23 72.97$\pm$1.47 28.38$\pm$4.23 52.76$\pm$3.41
DeWave[19] - 43.31$\pm$5.80 57.12$\pm$3.01 3.70$\pm$5.35 33.52$\pm$3.98
BrainBERT[49] 5.30$\pm$0.90 23.49$\pm$1.29 2.74$\pm$0.40 5.18$\pm$0.38
@@
Du-IN (poms) 73.23$\pm$0.67 79.12$\pm$1.07 28.89$\pm$1.96 64.24$\pm$1.29
  • 1
    PT: Whether the model is pre-trained before evaluation.
  • 2
    MS: Whether the model is pre-trained across multiple subjects.
@@
Methods Config Accuracy (%) $\pm$ Std (%)
PT11footnotemark: 1 MS22footnotemark: 2 subj-05 subj-06 subj-07 subj-08
TS-TCC[20] 34.53$\pm$1.34 9.73$\pm$0.97 24.83$\pm$1.15 20.08$\pm$1.60
CNN-BiGRU[37] - 51.26$\pm$4.93 31.52$\pm$1.48 47.75$\pm$1.12 24.64$\pm$4.44
EEG-Conformer[44] - 65.44$\pm$1.31 31.06$\pm$2.58 47.89$\pm$1.86 42.12$\pm$2.08
Neuro-BERT[50] 71.61$\pm$1.97 36.63$\pm$2.15 50.23$\pm$2.63 43.24$\pm$1.36
DeWave[19] - 45.20$\pm$3.08 26.88$\pm$1.92 38.24$\pm$2.15 29.15$\pm$3.28
BrainBERT[49] 9.59$\pm$0.90 2.67$\pm$0.31 4.79$\pm$0.64 5.10$\pm$0.59
@@
Du-IN (poms) 76.91$\pm$1.38 47.01$\pm$1.76 61.68$\pm$0.95 55.17$\pm$1.37
  • 1
    PT: Whether the model is pre-trained before evaluation.
  • 2
    MS: Whether the model is pre-trained across multiple subjects.
@@
Methods Config Accuracy (%) $\pm$ Std (%)
PT11footnotemark: 1 MS22footnotemark: 2 subj-09 subj-10 subj-11 subj-12
TS-TCC[20] 37.75$\pm$1.22 5.71$\pm$0.38 35.72$\pm$0.95 14.12$\pm$0.68
CNN-BiGRU[37] - 44.03$\pm$5.88 7.11$\pm$0.71 28.44$\pm$3.42 13.17$\pm$3.41
EEG-Conformer[44] - 56.51$\pm$1.98 22.22$\pm$1.07 57.10$\pm$2.03 29.87$\pm$1.44
Neuro-BERT[50] 54.12$\pm$4.11 24.66$\pm$1.28 62.99$\pm$0.93 36.07$\pm$2.36
DeWave[19] - 41.98$\pm$4.60 6.22$\pm$0.94 44.60$\pm$3.35 19.22$\pm$2.95
BrainBERT[49] 6.82$\pm$1.42 2.55$\pm$0.49 8.73$\pm$0.92 3.71$\pm$1.02
@@
Du-IN (poms) 70.71$\pm$1.48 36.90$\pm$1.34 72.80$\pm$1.38 43.53$\pm$3.20
  • 1
    PT: Whether the model is pre-trained before evaluation.
  • 2
    MS: Whether the model is pre-trained across multiple subjects.
Appendix L Effectiveness of Vector-Quantized Neural Signal Prediction
To verify the effectiveness of vector-quantized neural signal prediction, we elaborate on two types of experimental settings as illustrated in Table 13. The comparison between Du-IN and Setting 1 demonstrates that the codex is effective for masked sEEG modeling. The comparison between Du-IN and Setting 2 demonstrates that introducing the codex can prevent the model from focusing too much on reconstructing details, thus enabling the Du-IN MAE to learn better contextual embeddings.
Table 13: Ablations to validate the effectiveness of vector-quantized neural signal prediction.
       Model        Du-IN (mae)        Setting 1        Setting 2
       Acc. (%) $\pm$ Ste (%)        62.70$\pm$4.69        60.92$\pm$4.38        58.72$\pm$5.02
  • 1
    Setting 1: We directly predict output embeddings of the Du-IN Encoder in the Du-IN VQ-VAE by maximizing cosine similarity instead of predicting the discrete neural tokens from the codex.
  • 2
    Setting 2: We discard the Du-IN Encoder in the Du-IN VQ-VAE and directly reconstruct raw EEG patches by minimizing MSE loss.
Appendix M Ablation on Mask Ratio
In this experiment, we conduct different mask ratio settings to explore its impact. It is noted that we introduce the symmetric masking strategy, so we only need to validate half of the mask ratios. As the mask ratio is set to $r$, the symmetric masking will mask $1-r$ proportion of sEEG patches. The ablation results are provided in Table 14. It can be induced that the best mask ratio is 0.5 (0.5) for our dataset.
Table 14: Ablations to explore the impact of mask ratios.
Mask Ratio 0.5 (0.5) 0.4 (0.6) 0.3 (0.7) 0.2 (0.8) 0.1 (0.9)
Acc. (%) $\pm$ Ste (%) 62.70$\pm$4.69 60.58$\pm$4.33 59.58$\pm$4.98 58.92$\pm$4.07 58.55$\pm$3.94
Appendix N Ablation on Pre-training Epochs
The impact of the number of pre-training epochs (of the Du-IN VQ-VAE model) is demonstrated in Table 15. We use the checkpoints according to the specified epochs to pre-train the Du-IN MAE model for 400 epochs. Once the reconstruction loss of the Du-IN VQ-VAE model converges, the Du-IN VQ-VAE model can extract the state of the brain region well, thus leading to better performance.
The impact of the number of pre-training epochs (of the Du-IN MAE model) is demonstrated in Table 16. We use the checkpoints according to the specified epochs for downstream classification. Once the mask modeling loss of the Du-IN MAE model converges, the Du-IN MAE model learns robust contextual embeddings, thus leading to better performance.
Table 15: Ablations to explore the impact of the pre-training epochs (of the Du-IN VQ-VAE model).
# of Epochs 5 10 50 100 400
Acc. (%) $\pm$ Ste (%) 50.02$\pm$4.91 52.29$\pm$5.09 61.09$\pm$4.28 62.59$\pm$4.32 62.70$\pm$4.69
Table 16: Ablations to explore the impact of the pre-training epochs (of the Du-IN MAE model).
# of Epochs 5 10 50 100 400
Acc. (%) $\pm$ Ste (%) 57.87$\pm$4.58 58.12$\pm$4.49 61.89$\pm$4.62 62.47$\pm$4.77 62.70$\pm$4.69
Appendix O Subject-Wise Electrode Locations
@@
Appendix P Subject-Wise Selected Channels
The MNI coordinates, and brain region labels (according to Harvard-Oxford cortical and subcortical structural atlases [15]) for selected channels are listed below. The channels for each subject are arranged in descending order based on their contribution scores.
Table 17: The MNI coordinates and brain region labels of selected channels from subjects (01-04).
Subjects MNI coordinate Brain Region Du-IN Accuracy (%)
x y z
subj-01 -57 -16 19 Central Opercular Cortex L 71.25
-61 -16 20 Postcentral Gyrus L
-54 -16 17 Central Opercular Cortex L
-67 -15 23 Postcentral Gyrus L
-25 -30 49 White L
-64 -15 21 Postcentral Gyrus L
-51 -17 16 Central Opercular Cortex L
-22 -31 49 White L
-48 -17 15 Central Opercular Cortex L
-18 -32 49 White L
subj-02 33 -27 7 White R 77.99
37 -28 7 Heschls Gyrus R
30 -27 6 White R
48 -29 9 Planum temporale R
51 -29 10 Planum temporale R
41 -28 8 White R
46 -3 -9 Insula R
55 -29 11 Planum temporale R
49 -3 -7 Planum temporale R
43 -3 -10 Insula R
subj-03 -38 -30 6 White L 23.04
-58 -31 4 Superior Temporal Gyrus L
-31 -30 7 White L
-35 -30 7 White L
49 -11 1 Heschls Gyrus R
48 -36 26 Parietal Operculum Cortex R
42 -11 -1 Insula R
-55 -31 5 Superior Temporal Gyrus L
-41 -30 6 Planum temporale L
45 -11 0 Heschls Gyrus R
subj-04 -44 -10 32 Precentral Gyrus L 59.91
-45 -11 35 Precentral Gyrus L
-53 -6 -1 Planum temporale L
-44 -9 28 Precentral Gyrus L
-46 -12 39 Precentral Gyrus L
-52 -5 2 Planum temporale L
-43 -8 24 White L
-38 -3 6 Insula L
-53 -7 -5 Superior Temporal Gyrus L
-16 43 25 White L
Table 18: The MNI coordinates and brain region labels of selected channels from subjects (05-08).
Subjects MNI coordinate Brain Region Du-IN Accuracy (%)
x y z
subj-05 -48 -16 33 Postcentral Gyrus L 77.60
-46 15 29 Postcentral Gyrus L
-43 -14 22 White L
-51 -17 39 Postcentral Gyrus L
-44 -15 26 White L
-53 -17 43 Postcentral Gyrus L
-50 -16 36 Postcentral Gyrus L
-41 -14 19 Insula L
-55 -18 46 Postcentral Gyrus L
-49 -37 9 White L
subj-06 56 -1 10 Central Opercular Cortex R 41.91
58 -4 4 Planum temporale R
52 4 21 Precentral Gyrus R
53 3 17 Precentral Gyrus R
57 -2 7 Central Opercular Cortex R
54 1 14 Precentral Gyrus R
51 6 24 Precentral Gyrus R
64 -25 -4 Middle Temporal Gyrus R
21 -1 11 White R
49 7 28 Precentral Gyrus R
subj-07 -38 -18 2 Insula L 59.63
-44 -23 1 Heschls Gyrus L
-41 -21 1 Heschls Gyrus L
-35 -16 2 Insula L
-50 -28 0 Superior Temporal Gyrus L
-47 -26 1 Planum temporale L
-52 -30 0 Superior Temporal Gyrus L
-26 -16 40 White L
-42 0 -8 Insula L
-39 -22 41 Postcentral Gyrus L
subj-08 -40 -20 4 Heschls Gyrus L 52.35
-43 -21 4 Heschls Gyrus L
-37 -20 4 Insula L
-50 -22 4 Heschls Gyrus L
-47 -21 4 Heschls Gyrus L
-55 3 24 Precentral Gyrus L
-64 -24 3 Superior Temporal Gyrus L
-61 -23 3 Superior Temporal Gyrus L
-54 -22 3 Planum temporale L
-52 2 22 Planum temporale L
Table 19: The MNI coordinates and brain region labels of selected channels from subjects (09-12).
Subjects MNI coordinate Brain Region Du-IN Accuracy (%)
x y z
subj-09 -58 -13 21 Postcentral Gyrus L 66.39
-55 -12 19 Central Opercular Cortex L
-53 -11 17 Central Opercular Cortex L
-50 -10 15 Central Opercular Cortex L
-61 -15 23 Postcentral Gyrus L
-63 -16 25 Postcentral Gyrus L
-45 -8 10 Central Opercular Cortex L
-47 -9 12 Central Opercular Cortex L
-42 -7 8 Insula L
52 -2 26 Precentral Gyrus R
subj-10 -34 -47 41 Supramarginal Gyrus L 27.07
-42 -55 41 Angular Gyrus L
-39 -53 41 Angular Gyrus L
-37 -50 41 Supramarginal Gyrus L
-25 -37 42 White L
-44 -58 41 Angular Gyrus L
-27 -39 42 White L
-18 -41 48 Postcentral Gyrus L
-34 -34 -23 Temporal Fusiform Cortex L
-13 -40 43 Precuneous Cortex L
subj-11 39 -22 3 Heschls Gyrus R 73.56
36 -23 3 Insula R
47 -22 2 White R
32 -23 4 White R
43 -22 2 Planum temporale R
53 -9 16 Central Opercular Cortex R
57 -8 17 Postcentral Gyrus R
50 -10 16 Central Opercular Cortex R
61 -21 0 Superior Temporal Gyrus R
39 -14 13 Insula R
subj-12 45 -20 11 Heschls Gyrus R 44.76
38 -20 12 Insula R
42 -20 11 Heschls Gyrus R
49 -20 10 Heschls Gyrus R
53 -19 10 Heschls Gyrus R
60 -19 9 Planum temporale R
56 -19 9 Planum temporale R
60 -58 3 Middle Temporal Gyrus R
56 -57 3 Middle Temporal Gyrus R
35 -21 12 Insula R